Prove that the length of the chord of a parabola which passes through the focus and which is inclined at 30° to the axis of the parabola is four times the length of the latus rectum.


1
Expert's answer
2022-02-08T11:36:56-0500

Solution:

Equation of parabola in polar form is

"\begin{aligned}

&\frac{2 \mathrm{a}}{\mathrm{SP}}=1-\cos 30^{\circ} \\

&\Rightarrow \mathrm{SP}=\frac{2 \mathrm{a}}{1-\cos 30^{\circ}}

\end{aligned}"

"\frac{2 \mathrm{a}}{\mathrm{r}}=1-\cos \theta"


Let "\mathrm{PP}^{\prime}"  be the focal chord then parametric angles of P and P' are "30^{\circ}" and "\pi+ 30^{\circ}"  respectively then


Also 

"\frac{2 \mathrm{a}}{\mathrm{SP}^{\prime}}=1-\cos \left(\pi+30^{\circ}\right) 


\\\Rightarrow \mathrm{SP}^{\prime}=\frac{2 \mathrm{a}}{1-\cos \left(\pi+30^{\circ}\right)}"


Length of focal chord = PS + PS'

"\begin{aligned}

&\Rightarrow 1=\frac{2 \mathrm{a}}{1-\cos 30^{\circ}}+\frac{2 \mathrm{a}}{1-\cos \left(\pi+30^{\circ}\right)} \\

&\Rightarrow 1=\frac{2 \mathrm{a}}{1-\cos 30^{\circ}}+\frac{2 \mathrm{a}}{1+\cos 30^{\circ}}

\end{aligned}"

"\begin{aligned}
&\Rightarrow 1=2 \mathrm{a}\left\{\frac{1+\cos 30^{\circ}+1-\cos 30^{\circ}}{1-\cos ^{2} 30^{\circ}}\right\} \\
&\Rightarrow 1=2 \mathrm{a}\left\{\frac{2}{1-\frac{3}{4}}\right\} \\
&\Rightarrow 1=16 \mathrm{a} \\
&\Rightarrow 1=4 \times 4 \mathrm{a} \\
&\Rightarrow 1=4 \times \text { length of latus rectum } \\
&\text { Hence proved. }
\end{aligned}"